# Straggler-Resilient Decentralized Learning via Adaptive Asynchronous Updates

### v1

@@ C.1 Proof of Theorem 1
$\leq\mathbb{E}[F(\bar{{\mathbf{w}}}(k))]-\eta\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}+\frac{\eta}{2}\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}+\frac{\eta^{2}L}{2N}\sigma^{2}_{L}$
$+\frac{\eta^{2}L}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g({\mathbf{w}}_{j}(k))\right\|^{2}\!\!-\!\frac{\eta}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g({\mathbf{w}}_{j}(k))\right\|^{2}\!\!+\!\frac{\eta L^{2}}{2N}\sum\limits_{j=1}^{N}\mathbb{E}\left\|{\mathbf{w}}_{j}(k)\!-\!\bar{{\mathbf{w}}}(k)\right\|^{2}$
$\leq\mathbb{E}[F(\bar{{\mathbf{w}}}(k))]-\frac{\eta}{2}\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}+\frac{\eta^{2}L-\eta}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g({\mathbf{w}}_{j}(k))\right\|^{2}$
$+\frac{\eta L^{2}}{2N}\sum\limits_{j=1}^{N}\mathbb{E}\left\|{\mathbf{w}}_{j}(k)-\bar{{\mathbf{w}}}(k)\right\|^{2}+\frac{\eta^{2}L}{2N}\sigma_{L}^{2}.$ (14)
@@
$\qquad\qquad\qquad-\sum\limits_{i=1}^{N}{\mathbf{w}}_{i}(0)\Phi_{k:1}(j,i)+\eta\sum\limits_{s=1}^{k}\sum\limits_{i=1}^{N}g_{i}({\mathbf{w}}_{i}(s-1))\Phi_{k:s}(j,i)\Bigg{\|}^{2}$
$=\mathbb{E}\Bigg{\|}\sum\limits_{i=1}^{N}{\mathbf{w}}_{i}(0)(\frac{1}{N}-\Phi_{k:1}(j,i))-\eta\sum\limits_{s=1}^{k}\sum\limits_{i=1}^{N}g_{i}({\mathbf{w}}_{i}(s-1))(\frac{1}{N}-\Phi_{k:s}(j,i))\Bigg{\|}^{2}$
$\overset{(b_{1})}{\leq}\mathbb{E}\left\|\sum\limits_{i=1}^{N}{\mathbf{w}}_{i}(0)(\frac{1}{N}-\Phi_{k:1}(i,j))\right\|^{2}+\mathbb{E}\left\|\eta\sum\limits_{s=1}^{k}\sum\limits_{i=1}^{N}g_{i}({\mathbf{w}}_{i}(s-1))(\frac{1}{N}-\Phi_{k:s}(i,j))\right\|^{2}$
@@
$\qquad\qquad\qquad+\mathbb{E}\Bigg{\|}\eta\sum\limits_{s=1}^{k}\sum\limits_{i=1}^{N}2g({\mathbf{w}}_{i}(s-1))\frac{1+\beta^{-NB}}{1-\beta^{NB}}(1-\beta^{NB})^{(k-s)/NB}\Bigg{\|}^{2}$
@@
$+\!\mathbb{E}4\eta^{2}C^{2}\sum\limits_{s=1}^{k}\sum\limits_{s^{\prime}=1}^{k}q^{(2k-s-s^{\prime})}\Bigg{\|}\sum\limits_{i=1}^{N}(g_{i}({\mathbf{w}}_{i}(s-1))\!\!-\!\!\nabla F_{i}({\mathbf{w}}_{i}(s-1)))\Bigg{\|}\!\cdot\!\Bigg{\|}\sum\limits_{i=1}^{N}\nabla F_{i}({\mathbf{w}}_{i}(s-1))\Bigg{\|}$
$\qquad\qquad\qquad+2\mathbb{E}\Bigg{\|}\eta C\sum\limits_{s=1}^{k}q^{(k-s)}\sum\limits_{i=1}^{N}\nabla F_{i}({\mathbf{w}}_{i}(s-1))\Bigg{\|}^{2}$
$\overset{(b_{5})}{\leq}\mathbb{E}\frac{10\eta^{2}C^{2}}{1-q}\sum_{s=1}^{k}q^{k-s}\Bigg{\|}\sum\limits_{i=1}^{N}(g_{i}({\mathbf{w}}_{i}(s-1))-\nabla F_{i}({\mathbf{w}}_{i}(s-1)))\Bigg{\|}^{2}\allowdisplaybreaks$
@@
$\overset{(b_{6})}{\leq}\frac{10\eta^{2}C^{2}N^{2}}{(1-q)^{2}}\sigma^{2}+\frac{30\eta^{2}C^{2}N^{2}}{(1-q)^{2}}\varsigma^{2}+\frac{30\eta^{2}C^{2}L^{2}}{1-q}\sum_{s=1}^{k}q^{k-s}\sum_{i=1}^{N}\mathbb{E}\|{\mathbf{w}}_{i}(s)-{\mathbf{y}}(s)\|^{2}$
$\qquad\qquad\qquad+\frac{30\eta^{2}C^{2}N^{2}}{1-q}\sum_{s=1}^{k}q^{k-s}\mathbb{E}\|\nabla F({\mathbf{y}}(s))\|^{2},$ (15)
where the inequality $(b_{1})$ is due to the inequality $\|{\mathbf{a}}-{\mathbf{b}}\|^{2}\leq\|{\mathbf{a}}\|^{2}+\|{\mathbf{b}}\|^{2}$. $(b_{2})$ holds according to Lemma 2 (see the “Auxiliary Lemmas” Section below). W.l.o.g., we assume the initial term ${\mathbf{w}}_{i}(0),\forall i$ is small enough and we can neglect it. $(b_{3})$ follows $C:=2\cdot\frac{1+\beta^{-NB}}{1-\beta^{NB}}$ and $q:=(1-\beta^{NB})^{1/NB}$. $(b_{4})$ is due to $\|{\mathbf{a}}+{\mathbf{b}}\|^{2}\leq\|{\mathbf{a}}\|^{2}+\|{\mathbf{b}}\|^{2}+2{\mathbf{a}}{\mathbf{b}}$ and $(b_{5})$ is some standard mathematical manipulation by leveraging the following inequality, i.e., for any $q\in(0,1)$ and non-negative sequence $\{\chi(s)\}_{s=0}^{k}$, it holds [assran2019stochastic] $\sum_{k=0}^{K}\sum_{s=0}^{k}q^{k-s}\chi(s)\leq\frac{1}{1-\lambda}\sum_{s=0}^{K}\chi(s).$ $(b_{6})$ holds due to Assumption 4.4, the inequality
@@
$\qquad\qquad\qquad+3\mathbb{E}\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}$
@@
Rearranging the order of each term in (C.1) and letting $L\eta\leq 1$, we have
$\frac{\eta}{2}\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}$ $\leq\mathbb{E}[F(\bar{{\mathbf{w}}}(k))]-\mathbb{E}[F(\bar{{\mathbf{w}}}(k+1))]+\frac{\eta L^{2}}{2N}\sum\limits_{j=1}^{N}\mathbb{E}\left\|{\mathbf{w}}_{j}(k)-\bar{{\mathbf{w}}}(k)\right\|^{2}$
$\qquad\qquad\qquad+\frac{L\eta^{2}}{2N}\sigma^{2}_{L}+\frac{\eta}{2}\sigma_{L}^{2}+\underset{\leq 0}{\underbrace{\frac{\eta^{2}L-\eta}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g({\mathbf{w}}_{j}(k))\right\|^{2}}}$
@@
$\qquad\qquad\qquad+\frac{\eta L^{2}}{2N}\sum\limits_{j=1}^{N}\mathbb{E}\left\|{\mathbf{w}}_{j}(k)-\bar{{\mathbf{w}}}(k)\right\|^{2}.$ (16)
@@
$\qquad\qquad+\frac{30\eta^{2}C^{2}N^{3}}{1-q}\sum_{s=1}^{k}q^{k-s}\mathbb{E}\|\nabla F(\bar{{\mathbf{w}}}(s))\|^{2}+\frac{10\eta^{2}C^{2}N^{3}}{(1-q)^{2}}\sigma^{2}+\frac{30\eta^{2}C^{2}N^{2}}{(1-q)^{2}}\varsigma^{2}\Bigg{)}$
@@
$\qquad\qquad+\frac{30\eta^{2}C^{2}N^{3}}{(1-q)^{2}}\sum_{k=0}^{K-1}\mathbb{E}\|\nabla F(\bar{{\mathbf{w}}}(s))\|^{2}+\frac{10K\eta^{2}C^{2}N^{3}}{(1-q)^{2}}\sigma^{2}+\frac{30K\eta^{2}C^{2}N^{2}}{(1-q)^{2}}\varsigma^{2}\Bigg{)}$
$\leq\frac{30\eta^{2}C^{2}N^{3}}{(1-q)^{2}-30N\eta^{2}C^{2}L^{2}}\sum_{k=0}^{K-1}\mathbb{E}\|\nabla F(\bar{{\mathbf{w}}}(s))\|^{2}+\frac{10K\eta^{2}C^{2}N^{3}}{(1-q)^{2}-30N\eta^{2}C^{2}L^{2}}\sigma^{2}$
$\qquad\qquad\qquad+\frac{30K\eta^{2}C^{2}N^{3}}{(1-q)^{2}-30N\eta^{2}C^{2}L^{2}}\varsigma^{2},$
summing the recursion in (C.1) from the 00-th iteration to the $(K-1)$-th iteration yields
$\sum\limits_{k=0}^{K-1}\frac{\eta}{2}\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}$ $\leq\mathbb{E}[F(\bar{{\mathbf{w}}}(0))]-\mathbb{E}[F(\bar{{\mathbf{w}}}(K))]+\frac{KL\eta^{2}}{2N}\sigma^{2}_{L}\allowdisplaybreaks$
@@
$\qquad+\frac{5K\eta^{3}C^{2}N^{2}L^{2}}{(1-q)^{2}-30N\eta^{2}C^{2}L^{2}}\sigma^{2}+\frac{15K\eta^{3}C^{2}N^{2}L^{2}}{(1-q)^{2}-30N\eta^{2}C^{2}L^{2}}\varsigma^{2}.$ (17)
@@
$\qquad\leq\mathbb{E}[F(\bar{{\mathbf{w}}}(0))]-\mathbb{E}[F(\bar{{\mathbf{w}}}(K))]+\frac{KL\eta^{2}}{2N}\sigma^{2}_{L}$
$\qquad+\frac{5K\eta^{3}C^{2}N^{2}L^{2}}{(1-q)^{2}-30N\eta^{2}C^{2}L^{2}}\sigma_{L}^{2}+\frac{15K\eta^{3}C^{2}N^{2}L^{2}}{(1-q)^{2}-30N\eta^{2}C^{2}L^{2}}\varsigma^{2}.$ (18)
@@
$\frac{\eta}{6}\sum\limits_{k=0}^{K-1}$ $\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}\leq\mathbb{E}[F(\bar{{\mathbf{w}}}(0))]-\mathbb{E}[F(\bar{{\mathbf{w}}}(K))]+\frac{KL\eta^{2}}{2N}\sigma^{2}_{L}+\frac{K\eta^{2}}{9N}\sigma_{L}^{2}+\frac{K\eta^{2}}{3N}\varsigma^{2}.$ (19)
@@
$\frac{1}{K}\sum\limits_{k=0}^{K-1}\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}\leq\mathbb{E}\frac{6[F(\bar{{\mathbf{w}}}(0))]-\mathbb{E}[F(\bar{{\mathbf{w}}}(K))]}{\eta K}+\frac{9L\eta+2\eta}{3N}\sigma^{2}_{L}+\frac{2\eta}{N}\varsigma^{2}.$ (20)
@@
C.2 Proof of Corollary 1
Corollary 2.
@@ C.2 Proof of Corollary 1
The desired results is yield by substituting $\eta=\sqrt{N/K}$ back in to (7) in Theorem 1. ∎
C.3 Auxiliary Lemmas
We provide the following auxiliary lemmas which are used in our proofs. We omit the proofs of these lemmas for the ease of exposition and refer interested readers to [Boyd05] and [nedic2009distributed] for details.
Lemma 1 (Theorem 2 in [Boyd05]).

### v2

@@ Appendix B Proofs of Main Result
$\leq\mathbb{E}[F(\bar{{\mathbf{w}}}(k))]-\eta\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}+\frac{\eta}{2}\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}$
$+\!\frac{\eta^{2}L}{2N}\sigma^{2}_{L}\!+\!\frac{\eta^{2}L}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g({\mathbf{w}}_{j}(k))\right\|^{2}\!\!\!\!-\!\frac{\eta}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g({\mathbf{w}}_{j}(k))\right\|^{2}\!\!\allowdisplaybreaks$
$\qquad+\!\frac{\eta L^{2}}{2N}\sum\limits_{j=1}^{N}\mathbb{E}\left\|{\mathbf{w}}_{j}(k)\!-\!\bar{{\mathbf{w}}}(k)\right\|^{2}$
$\leq\mathbb{E}[F(\bar{{\mathbf{w}}}(k))]-\frac{\eta}{2}\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}+\frac{\eta^{2}L}{2N}\sigma_{L}^{2}$
(14) $+\!\frac{\eta^{2}L\!-\!\eta}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g({\mathbf{w}}_{j}(k))\right\|^{2}\!\!\!+\!\frac{\eta L^{2}}{2N}\!\sum\limits_{j=1}^{N}\!\mathbb{E}\!\left\|{\mathbf{w}}_{j}(k)\!-\!\bar{{\mathbf{w}}}(k)\right\|^{2}.$
@@
$-\sum\limits_{i=1}^{N}{\mathbf{w}}_{i}(0)\Phi_{k:1}(j,i)+\eta\sum\limits_{s=1}^{k}\sum\limits_{i=1}^{N}g_{i}({\mathbf{w}}_{i}(s-1))\Phi_{k:s}(j,i)\Bigg{\|}^{2}$
$=\mathbb{E}\Bigg{\|}\sum\limits_{i=1}^{N}{\mathbf{w}}_{i}(0)(\frac{1}{N}-\Phi_{k:1}(j,i))\allowdisplaybreaks$
$\qquad-\eta\sum\limits_{s=1}^{k}\sum\limits_{i=1}^{N}g_{i}({\mathbf{w}}_{i}(s-1))(\frac{1}{N}-\Phi_{k:s}(j,i))\Bigg{\|}^{2}$
$\overset{(b_{1})}{\leq}\mathbb{E}\left\|\sum\limits_{i=1}^{N}{\mathbf{w}}_{i}(0)(\frac{1}{N}-\Phi_{k:1}(i,j))\right\|^{2}\allowdisplaybreaks$
$\qquad\qquad+\mathbb{E}\left\|\eta\sum\limits_{s=1}^{k}\sum\limits_{i=1}^{N}g_{i}({\mathbf{w}}_{i}(s-1))(\frac{1}{N}-\Phi_{k:s}(i,j))\right\|^{2}$
@@
$+\!\!\mathbb{E}\Bigg{\|}\eta\sum\limits_{s=1}^{k}\!\sum\limits_{i=1}^{N}\!2g({\mathbf{w}}_{i}(s\!-\!1))\frac{1\!+\!\beta^{-NB}}{1\!-\!\beta^{NB}}(1\!-\!\beta^{NB})^{(k\!-\!s)/NB}\Bigg{\|}^{2}$
@@
$+\!\mathbb{E}4\eta^{2}C^{2}\sum\limits_{s=1}^{k}\sum\limits_{s^{\prime}=1}^{k}q^{(2k-s-s^{\prime})}$
$\Bigg{\|}\!\sum\limits_{i=1}^{N}\!(g_{i}({\mathbf{w}}_{i}(s\!-\!1))\!\!-\!\!\nabla F_{i}({\mathbf{w}}_{i}(s\!-\!1)))\Bigg{\|}\!\cdot\!\!\Bigg{\|}\sum\limits_{i=1}^{N}\!\nabla F_{i}({\mathbf{w}}_{i}(s\!-\!1))\Bigg{\|}$
$\qquad\qquad+2\mathbb{E}\Bigg{\|}\eta C\sum\limits_{s=1}^{k}q^{(k-s)}\sum\limits_{i=1}^{N}\nabla F_{i}({\mathbf{w}}_{i}(s-1))\Bigg{\|}^{2}$
$\overset{(b_{5})}{\leq}\!\!\mathbb{E}\frac{10\eta^{2}C^{2}}{1-q}\!\!\sum_{s=1}^{k}\!q^{k-s}\Bigg{\|}\!\sum\limits_{i=1}^{N}(g_{i}({\mathbf{w}}_{i}(s\!-\!1))\!\!-\!\!\nabla F_{i}({\mathbf{w}}_{i}(s\!-\!1)))\Bigg{\|}^{2}$
@@
$\overset{(b_{6})}{\leq}\frac{10\eta^{2}C^{2}N^{2}}{(1-q)^{2}}\sigma^{2}+\frac{30\eta^{2}C^{2}N^{2}}{(1-q)^{2}}\varsigma^{2}$
$\qquad+\frac{30\eta^{2}C^{2}L^{2}}{1-q}\sum_{s=1}^{k}q^{k-s}\sum_{i=1}^{N}\mathbb{E}\|{\mathbf{w}}_{i}(s)-{\mathbf{y}}(s)\|^{2}$
(15) $\qquad\qquad\qquad+\frac{30\eta^{2}C^{2}N^{2}}{1-q}\sum_{s=1}^{k}q^{k-s}\mathbb{E}\|\nabla F({\mathbf{y}}(s))\|^{2},$
where the inequality $(b_{1})$ is due to the inequality $\|{\mathbf{a}}-{\mathbf{b}}\|^{2}\leq\|{\mathbf{a}}\|^{2}+\|{\mathbf{b}}\|^{2}$. $(b_{2})$ holds according to Lemma B.3 (see the “Auxiliary Lemmas” Section below). W.l.o.g., we assume the initial term ${\mathbf{w}}_{i}(0),\forall i$ is small enough so as to neglect it. $(b_{3})$ follows $C:=2\cdot\frac{1+\beta^{-NB}}{1-\beta^{NB}}$ and $q:=(1-\beta^{NB})^{1/NB}$. $(b_{4})$ is due to $\|{\mathbf{a}}+{\mathbf{b}}\|^{2}\leq\|{\mathbf{a}}\|^{2}+\|{\mathbf{b}}\|^{2}+2{\mathbf{a}}{\mathbf{b}}$ and $(b_{5})$ is some standard mathematical manipulation by leveraging the following inequality, i.e., for any $q\in(0,1)$ and non-negative sequence $\{\chi(s)\}_{s=0}^{k}$, it holds (Assran et al., 2019) $\sum_{k=0}^{K}\sum_{s=0}^{k}q^{k-s}\chi(s)\leq\frac{1}{1-\lambda}\sum_{s=0}^{K}\chi(s).$ $(b_{6})$ holds due to Assumption 4.4, the inequality
@@
[…]
@@
Rearranging the order of each term in (B.1) and letting $L\eta\leq 1$,
$\frac{\eta}{2}\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}$ $\leq\mathbb{E}[F(\bar{{\mathbf{w}}}(k))]-\mathbb{E}[F(\bar{{\mathbf{w}}}(k+1))]\allowdisplaybreaks$
$+\frac{\eta L^{2}}{2N}\sum\limits_{j=1}^{N}\mathbb{E}\left\|{\mathbf{w}}_{j}(k)-\bar{{\mathbf{w}}}(k)\right\|^{2}$
$+\frac{L\eta^{2}}{2N}\sigma^{2}_{L}+\frac{\eta}{2}\sigma_{L}^{2}+\underset{\leq 0}{\underbrace{\frac{\eta^{2}L-\eta}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g({\mathbf{w}}_{j}(k))\right\|^{2}}}$
@@
(16) $\qquad\qquad+\frac{\eta L^{2}}{2N}\sum\limits_{j=1}^{N}\mathbb{E}\left\|{\mathbf{w}}_{j}(k)-\bar{{\mathbf{w}}}(k)\right\|^{2}.$
@@
$+\frac{30\eta^{2}C^{2}N^{3}}{1-q}\sum_{s=1}^{k}q^{k-s}\mathbb{E}\|\nabla F(\bar{{\mathbf{w}}}(s))\|^{2}\allowdisplaybreaks$
$\qquad+\frac{10\eta^{2}C^{2}N^{3}}{(1-q)^{2}}\sigma^{2}+\frac{30\eta^{2}C^{2}N^{2}}{(1-q)^{2}}\varsigma^{2}\Bigg{)}$
@@
$\qquad+\frac{30\eta^{2}C^{2}N^{3}}{(1-q)^{2}}\sum_{k=0}^{K-1}\mathbb{E}\|\nabla F(\bar{{\mathbf{w}}}(s))\|^{2}\allowdisplaybreaks$
$\qquad+\frac{10K\eta^{2}C^{2}N^{3}}{(1-q)^{2}}\sigma^{2}+\frac{30K\eta^{2}C^{2}N^{2}}{(1-q)^{2}}\varsigma^{2}\Bigg{)}$
$\leq\frac{30\eta^{2}C^{2}N^{3}}{(1-q)^{2}-30N\eta^{2}C^{2}L^{2}}\sum_{k=0}^{K-1}\mathbb{E}\|\nabla F(\bar{{\mathbf{w}}}(s))\|^{2}\allowdisplaybreaks$
$+\frac{10K\eta^{2}C^{2}N^{3}}{(1-q)^{2}-30N\eta^{2}C^{2}L^{2}}\sigma^{2}+\frac{30K\eta^{2}C^{2}N^{3}}{(1-q)^{2}-30N\eta^{2}C^{2}L^{2}}\varsigma^{2},$
summing the recursion in (B.1) from the 00-th iteration to the $(K-1)$-th iteration yields
$\sum\limits_{k=0}^{K-1}\frac{\eta}{2}\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}\allowdisplaybreaks$
$\leq\mathbb{E}[F(\bar{{\mathbf{w}}}(0))]-\mathbb{E}[F(\bar{{\mathbf{w}}}(K))]+\frac{KL\eta^{2}}{2N}\sigma^{2}_{L}\allowdisplaybreaks$
@@
(17) $+\!\frac{5K\eta^{3}C^{2}N^{2}L^{2}}{(1\!\!-\!q)^{2}\!\!-\!30N\eta^{2}C^{2}L^{2}}\sigma^{2}\!+\!\frac{15K\eta^{3}C^{2}N^{2}L^{2}}{(1\!-\!q)^{2}\!-\!30N\eta^{2}C^{2}L^{2}}\varsigma^{2}.$
@@
(18) $+\frac{5K\eta^{3}C^{2}N^{2}L^{2}}{(1\!-\!q)^{2}\!-\!30N\eta^{2}C^{2}L^{2}}\sigma_{L}^{2}\!+\!\frac{15K\eta^{3}C^{2}N^{2}L^{2}}{(1-q)^{2}\!-\!30N\eta^{2}C^{2}L^{2}}\varsigma^{2}.$
@@
$\frac{\eta}{6}\sum\limits_{k=0}^{K-1}$ $\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}\leq\mathbb{E}[F(\bar{{\mathbf{w}}}(0))]-\mathbb{E}[F(\bar{{\mathbf{w}}}(K))]\allowdisplaybreaks$
(19) $+\frac{KL\eta^{2}}{2N}\sigma^{2}_{L}+\frac{K\eta^{2}}{9N}\sigma_{L}^{2}+\frac{K\eta^{2}}{3N}\varsigma^{2}.$
@@
$\frac{1}{K}\sum\limits_{k=0}^{K-1}\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}\leq\mathbb{E}\frac{6[F(\bar{{\mathbf{w}}}(0))]-\mathbb{E}[F(\bar{{\mathbf{w}}}(K))]}{\eta K}\allowdisplaybreaks$
(20) $\qquad\qquad+\frac{9L\eta+2\eta}{3N}\sigma^{2}_{L}+\frac{2\eta}{N}\varsigma^{2}.$
@@
B.2. Proof of Corollary 4.2
Corollary B.1.
@@
The desired results is yield by substituting $\eta=\sqrt{N/K}$ back in to (7) in Theorem 4.1. ∎
B.3. Auxiliary Lemmas
We provide the following auxiliary lemmas which are used in our proofs. We omit the proofs of these lemmas for ease of exposition and refer interested readers to (Xiao et al., 2006) and (Nedic and Ozdaglar, 2009).
Lemma B.2 (Theorem 2 in (Xiao et al., 2006)).